Providence Union clinic faces a monthly demand equation of:

Visits= 400-3*Price

  1. How many visits does the clinic need to have to be able to charge $50 per visit?
  2. What Price Should they charge if they are only able to have 100 visits per month?
  3. True or False: cutting Price from $80 to $70 will increase sales and revenue
  4. What type of elasticity does this clinic face?
  5. Calculate the price elasticity of demand for this clinic
Answer #1

Visits = 400 - 3 Price P = So visits = 400 - 3 (50) = 250 Visits = 100 10o = you -3P P = loo P = 80 visits = 400 - 3(80) = 16